Robert has 39 coins in his pocket. They are all dimes and quarters and total $6.00. How many of each does he have?
Answer by mananth(12270)

let dimes be x
quarters will be 39-x
value 600 cents
..
10x+25(39-x)=600
10x+975-25x=600
-15x=600-975
-15x=-375
x=-375/-15
x=25 dimes
14 will be quarters.
CHECK
25*14 +25*10= 600